OpenStudy (anonymous):

How do you simplify a radical that contains a perfect square factor? ((I searched Google and didn't get any results for some reason))

hartnn (hartnn):

like \(\sqrt{x^5}\) u have your own examples ? specific questions ?

OpenStudy (anonymous):

No specific questions; just wondering how to simplify /any/ radical which contains a perfect square factor ^-^

hartnn (hartnn):

ok, x^5 has a perfect square, which is ... ? can you find ?

OpenStudy (anonymous):

I honestly have no clue; I'm going to assume the perfect square is...0?... Yeah, I have no idea what I'm doing...

hartnn (hartnn):

OK, is x^4 a perfect square ? of whose ? or do you want to do this with numbers ? like \(\sqrt{128}\)

OpenStudy (anonymous):

128 doesn't have a perfect square, right? The closest I could find was 11.3137085

hartnn (hartnn):

128 is not a perfect square, but it has a perfect square in it like 27 27 = 3*3*3 = 9*3 9 is a perfect square! of 3 so, \(\sqrt{27}=\sqrt{3^2 \times 3}=3\times \sqrt3\) got this ? ask doubts...

OpenStudy (anonymous):

Oh, I see now

hartnn (hartnn):

similarly, \(\sqrt{128}= \sqrt{2*2*2*2*2*2*2}=\sqrt{8^2\times 2}=8\sqrt2\)